# Deep Graph Anomaly Detection: A Survey and New Perspectives

### v1

@@ 8 Conclusion
In this survey, we first discuss the complexities and existing challenges in GAD. Then we present a novel taxonomy for deep GAD methods from three new perspectives, including GNN backbone design, proxy task, and graph anomaly measures. We further deepen the discussions in each perspective by discussing more fine-grained categories of methods there. Along with each fine-grained methodology category, we not only review the associated GAD methods, but also analyze their general assumption, pros and cons, and their capabilities in addressing the unique challenges in GAD. We lastly discuss six important directions for future research on GAD. By tackling the problems in these directions, we expect much more advanced generation of methods for solving real-life GAD problems.
@@ Appendix A ALGORITHMS
Netwalk [140] Unsupervised Dynamic Edge Genuine Proxy Task Design 2018 Link
FraduNE [153] Supervised Static SubGraph Genuine Anomaly Measures 2018 Link
SemiGNN [119] Supervised Static Node Injected GNN Backbone 2019 N/A
AddGraph [152] Unsupervised Dynamic Edge Genuine Proxy Task Design 2019 Link
DOMINANT [18] Unsupervised Static Node Injected Proxy Task Design 2019 Link
GraphUCB[19] Unsupervised Static Node Injected Proxy Task Design 2019 Link
GraphConsis [73] Supervised Static Node Genuine GNN Backbone 2020 Link
CARE-GNN [25] Supervised Static Node Genuine GNN Backbone 2020 Link
AnomalyDAE[31] Unsupervised Static Node Injected Proxy Task Design 2020 Link
@@ Appendix C QUANTITATIVE COMPARISON
The way we compare the performance of different methods is to collect experimental results using the same dataset from their original papers. In this subsection, we do such an empirical comparison. Tables V and VI highlight the performance of several methods on both real and synthetic GAD datasets.
@@
Metric Method Dataset

### v2

@@ VIII Conclusion
In this survey, we first discuss the complexities and existing challenges in GAD. Then we present a novel taxonomy for deep GAD methods from three new perspectives, including GNN backbone design, proxy task, and graph anomaly measures. We further deepen the discussions in each perspective by discussing more fine-grained categories of methods. Along with each fine-grained methodology category, we not only review the associated GAD methods, but also analyze their general assumption, pros and cons, and their capabilities in addressing the unique challenges in GAD. We lastly discuss six important directions for future research on GAD. By tackling the problems in these six directions, we expect a more advanced generation of methods for solving real-life GAD tasks.
Acknowledgments
This research is supported in part by the National Research Foundation, Singapore and CyberSG R&D Programme Office under its Translation and Innovation Grants (CRPO-GC4-SMU-001), the Ministry of Education, Singapore under its Tier 1 Academic Research Fund (24-SIS-SMU-008), and the Lee Kong Chian Fellowship (T050273).
@@ Appendix A ALGORITHMS
Netwalk [142] UnsuperviWd Static Node Injected Proxy Task Design 2020 Link
@@ Appendix C QUANTITATIVE COMPARISON
In this subsection, we perform an empirical comparison of different graph anomaly detection methods. The way we compare the performance of different methods is to collect experimental results using the same datasets from their original papers. Tables V and VI provide the results of AUROC and AUPRC, two widely used metrics in anomaly detection, for various methods on both synthetic and real datasets for node-level anomaly detection, while Table VII presents the results of graph-level anomaly detectors on both homogeneous and heterogeneous graph datasets. Additionally, we collect the F1 performance under two setting with 40% and 1% labeled nodes used during training on four datasets, with the results shown in Table VIII. To evaluate the scalability of each method, we also provide a comparison of their runtime in Table IX. It should be noted that the runtime may not be directly comparable across the methods, as these computational costs may be obtained using different experimental environments.
Appendix D EXPLANATION/ANALYSIS ON PERFORMANCE COMPARISON
From the empirical comparison results, we can observe that supervised methods generally outperform unsupervised methods, as they leverage labeled information in their detection model learning. However, this does not always hold, particularly on datasets with injected anomalies, where unsupervised methods like reconstruction-based and contrastive learning-based methods achieve the best performance. This may be because supervised methods tend to overfit the limited training anomaly data, while the designs in some unsupervised methods are built upon some prior knowledge that is used for generating the injected anomalies. Additionally, the superior performance of most methods are limited to specific datasets due to their specialized design tailored to those datasets. Some methods such as reconstruction-based are unable to scale up to very large-scale datasets, such as T-Social and Dgraph. For graph-level anomaly detection, only AUROC results are typically available, as AUPRC is not reported in most studies in this line. Most graph-level anomaly detection methods are unsupervised and designed for homogeneous graphs.
Since AUROC and AUPRC are widely used metrics for anomaly detection tasks, only a few studies report F1 performance, and all of these methods are supervised. As the number of labeled nodes increases, the F1 performance generally gets improved.
For the runtime of each method, unsupervised methods typically require significantly more time than semi-supervised and supervised methods due to their inherently high complexity, often involving tasks such as reconstruction, contrastive learning, and local node affinity calculation. In contrast, supervised methods are optimized using a supervised loss function, which generally results in lower computational costs.
@@ Appendix D EXPLANATION/ANALYSIS ON PERFORMANCE COMPARISON
Metric Method Homogeneous Graphs Heterogeneous Graphs
@@
TABLE VIII: F1 comparison of node-level anomaly detection
Metric Method Datasets
Amazon YelpChi T-Finance T-Social
40% CARE-GNN [24] 0.863 0.633 0.825 0.518
PC-GNN [67] 0.895 0.630 0.558 0.441
BWGNN [117] 0.917 0.769 0.886 0.840
H2-FDetector [111] 0.869 0.743 0.742 /
GHRN [35] 0.923 0.775 0.879 0.682
GDN [93] 0.906 0.760 0.887 0.568
CONSISGAD [14] / / / /
PMP [167] 0.920 0.819 0.919 0.952
GLHAD [38] 0.932 0.823 0.923 0.942
LEX-GNN [50] 0.934 0.863 / /
1% CARE-GNN [24] 0.757 0.616 0.836 /
PC-GNN [67] 0.852 0.642 0.869 0.496
BWGNN [117] 0.904 0.612 0.869 0.763
H2-FDetector [111] 0.724 0.639 / /
GHRN [35] 0.863 0.613 0.8 0.712
GDN [93] 0.897 0.648 0.766 0.557
CONSISGAD [14] 0.900 0.697 0.909 0.780
PMP [167] 0.877 0.686 0.888 0.845
GLHAD [38] / / / /
LEX-GNN [50] 0.873 0.697 / /
TABLE IX: Runtime (in seconds) comparison of node-level anomaly detection
Method Datasets
Amazon YelpChi T-Finance T-Social Reddit Elliptic DGraph Tolokers Questions
DOMINANT [18] 1,592 / 10,721 / 125 1,119 / / /
AnomalyDAE [30] 1,656 / 18,560 / 161 8,296 / / /
OCGNN [124] 765 / 5,717 / 162 3,517 / / /
AEGIS [17] 1,121 / 15,258 / 166 5,638 / / /
GAAN [15] 1,678 / 12,120 / 94 1,866 / / /
TAM [101] 5,050 102,232 17,360 / 432 13,200 / 5,668 11,603
GGAD [102] 658 / 9,345 / 368 5,146 488 / /
GADNR [107] 2,048 5,046 14,255 / 692 12,568 / 861 2,795
SmoothGNN [22] 7 19 16 4,877 7 205 2,924 6 32
PREM [87] 1,267 308 266 / 73 2,149 / 74 409
CARE-GNN [24] / / 572 9,159 / / / / /
PC-GNN [67] / / 736 13,958 / / / / /
BWGNN [117] / / 31 2,707 / / / / /